# Plurals: A System for Guiding LLMs Via Simulated Social Ensembles

### v3

@@ Abstract.
Recent debates raised concerns that language models may favor certain viewpoints. But what if the solution is not to aim for a “view from nowhere” but rather to leverage different viewpoints? We introduce Plurals, a system and Python library for pluralistic AI deliberation. Plurals consists of Agents (LLMs, optionally with personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is a generator of simulated social ensembles. Plurals integrates with government datasets to create nationally representative personas, includes deliberation templates inspired by democratic deliberation theory, and allows users to customize both information-sharing structures and deliberation behavior within Structures. Six case studies demonstrate fidelity to theoretical constructs and efficacy. Three randomized experiments show simulated focus groups produced output resonant with an online sample of the relevant audiences (chosen over zero-shot generation in 75% of trials). Plurals is both a paradigm and a concrete system for pluralistic AI.
@@ 1. Introduction
As an alternative to “bias-free” models, we introduce a new pluralistic AI system (sorensen_roadmap_2024, ), Plurals, that can accomplish these tasks. It is a public-facing Python library (Figure 1 for system overview, Figure 2 for code snippets, see here [^1] https://github.com/josh-ashkinaze/plurals for library). Plurals consists of Agents (optionally integrated with government datasets for nationally representative personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is an end-to-end generator of customizable “simulated social ensembles”. We incorporate interaction templates inspired by democratic deliberation theory and integration with government datasets for nationally representative personas. For example, to create an Agent representing a male California resident, our system samples a statistically representative citizen from American National Election Studies, and then uses the citizen’s demographics and political stances as an LLM prompt. We draw on democratic deliberation theory, which emphasizes dialogue between different views (bachtiger_deliberative_2018, ; marti_pluralism_2017, ), as a blueprint. Our work builds on research in deliberation (bachtiger_deliberative_2018, ; morrell_listening_2018, ; brown_deliberation_2018, ; smith_mini-publics_2018, ; fraser_rethinking_1990, ; habermas_structural_1991, ; marti_pluralism_2017, ), pluralistic sociotechnical systems (argyle_out_2023, ; lee_webuildai_2019, ; gordon_jury_2022, ; zhang_policykit_2020, ), and multi-agent AI alignment approaches (vahidov_pluralistic_2004, ; irving_ai_2018, ; talebirad_multi-agent_2023, ; kim_can_2024, ; chan_chateval_2023, ). To our knowledge, Plurals is the first general-purpose toolkit for pluralistic, multi-agent interactions modeled after democratic deliberation.
@@ 1.1. Brief System Overview
Plurals allows users to create simulated social ensembles with Agents, Structures, and Moderators: Agents complete tasks within Structures, which define how information is shared between Agents. Moderators can summarize multi-agent communication. Each abstraction is highly customizable. Agents can use various LLMs and have system instructions set manually, through persona generation, or via American National Election Studies (ANES) integration. Structures vary in information-sharing, complexity, and randomness. For example, users can define custom networks of Agents in a few lines of code (Figure 2). The behavior of Agents within Structures (how they should combine information from other agents) can be tuned via combination instructions. Moderators can summarize multi-agent communication. Our package comes pre-populated with templates for personas, combination instructions, and moderators—drawing on democratic deliberation theory and prior work.
@@ 2.3. Multi-Agent Systems for AI Alignment
Like our system, several multi-agent systems are explicitly designed with the goal of alignment (irving_ai_2018, ; pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ). Broadly, these systems typically center interactions between agents or agent roles. For example, several projects have explored the role of AI alignment through debate (irving_ai_2018, ; khan_debating_2024, )). Khan et al. (khan_debating_2024, ) propose a system where a weaker model picks an answer after seeing stronger models debate, and this process of debate leads to more truthful answers for the weaker model. Other multi-agent systems center agent roles (pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ; zhu_role-based_2008, ; vahidov_pluralistic_2004, )—the idea being that agents playing distinct parts can aid human decision-makers (vahidov_pluralistic_2004, ).
To this body of research, we offer several contributions. More theoretically, our abstractions are specifically grounded in the theory and practice of deliberation. More practically, because our system has support for Agents, Structures, and Moderators, it effectively enables users to customize both information-sharing (as in AI debate literature) and Agent roles (as in the AI role literature). We extend the debate paradigm by allowing for arbitrary information structures. A back-and-forth debate is of course just one of many possible informational structures. Our system contributes to the role-based literature by integrating with ANES, enabling users to quickly draw up nationally representative roles. We also design around deliberation—the space in between roles and information-sharing. For example, users can ablate the role of an Agent (i.e: their system instructions) and the combination instructions of an Agent. Finally, Plurals is a fully functioning Python package and not a one-off study. Hence, Plurals can operate as shared infrastructure. It makes multi-agent systems faster to set up and more accessible for researchers.
@@ 3.2. Modularity
Apart from the practical utility, this separation between Agents and Structures aligns with well-established social science frameworks. This conceptualization is most explicitly articulated in Structuration Theory by Anthony Giddens (giddens_constitution_1986, ), which explores the interplay between “agents” and larger “structures” agents exist in. Giddens aimed to transcend theories of behavior that centered exclusively on either individuals or societal structures. Similar distinctions appear across disciplines: individuals and environments in development psychology (radke-yarrow_individual_1991, ), person and situation in social psychology (furr_persons_2021, ), individual and field in sociology (grenfell_field_2012, ), and agent and enviornment in artificial intelligence (russell_artificial_1995, ). By using Agents and Structures as core abstractions [^2] We chose the terms “Agents” and “Structures” based on AI terminology and definitional precision. We adopt conventional AI terminology, where agents refer to autonomous entities capable of perceiving and acting within an environment (russell_artificial_1995, ). We opted for “structures” instead of “environment” to more accurately reflect our system. Structures specifically define information-sharing patterns and interaction protocols between agents, describing a more bounded space than “environment.” Unlike “environment,” which often implies indeterminacy or randomness, “structure” connotes intentional arrangement. The Cambridge Dictionary defines structure as “the way in which the parts of a system or object are arranged or organized, or a system arranged in this way” (dictionary_structure_2024, ). This definition aligns with a user-tuned abstraction., we create a modularity that resonates with different disciplines.
@@ 4.1.1. Component Description
(b) In a moderated ensemble, nationally-representative Agents brainstorm ways to encourage recycling. Then a moderator with a persona inspired by divergent creativity literature (ashkinaze_how_2024, ) summarizes responses with custom combination instructions.
@@ 4.1.2. Implementation
Specific personas can be inputted by the user (e.g: “A graphic designer”) or drawn from American National Election Studies (ANES) [^4] Specifically, we are using the ANES pilot dataset from February 2024., as in Argyle et al. (argyle_out_2023, ). When using ANES, our system finds a real individual satisfying some criteria and then creates a persona based on the totality of this individual’s attributes. Sampling is always probability-weighted, so the probability of a citizen being simulated matches their national sample probability weight. Because ANES is nationally representative, the marginal distribution of Plurals-generated personas matches that of the general population. Code snippet 2(d) (top panel), shows initializing Agents based on specific criteria (e.g: California resident below the age of 40) using the query_str method, which searches ANES through a Pandas string [^5] For accessibility we have a helper function which prints a human-readable mapping of ANES variables., For convenience, we also support an ideology method (ideology=’liberal’) and initializing randomly selected ANES citizens (persona=’random’, 2(a)). The latter can be used to quickly draw up nationally-representative “citizen assemblies” (2(b)).
ANES is just one possible generator of data-driven personas, and in future iterations, we aim to provide additional persona-generation methods. We chose ANES as our initial dataset for the following reasons. First, it has been used in prior work—most notably, Argyle et al. (argyle_out_2023, ). Second, ANES has data on political ideologies, supporting the core motivation of this system—testing whether LLM outputs can be improved through pluralism. Third, ANES is updated more frequently than other nationally-representative datasets like the U.S census.
@@ Results
In the zero-shot condition, we set the system instructions of GPT-4o to “You are an expert copywriter for an ad agency” and the user prompt was “Come up with a specific product for a solar panel company that would resonate with conservatives. Be very specific. Answer in 50 words only.” In the Plurals condition, the Moderator had the same instructions. However, that Moderator oversaw an ensemble of 10 simulated ANES conservatives (initialized using our ideology persona method and anes persona template) who were asked what features they personally would want in a solar panel company. The Moderator then came up with a 50-word solar panel idea after exposure to these simulated discussions. For 15 trials, we generated a solar panel company idea with zero-shot and Plurals.
@@
5.5. Efficacy: Liberals preferred homeless shelter ideas from a simulated focus group of liberals over zero-shot
@@ 5.6. Moderation: Using Plurals for LLM Guardrails
We began with Abercrombie et al.’s (abercrombie_collaborative_2024, ) typology of AI, algorithmic, and automation harms. We selected two specific harms—environmental and physical harms. For each harm, we crafted three user prompts that would trigger concerns in one category but not the other (Appendix I), testing the Moderator’s ability to discriminate between tasks based on their specific value sets. We initialized Moderators with specific value sets using a CoT system prompt that incorporated Abercrombie et al.’s language around typology definitions (Appendix I), instructing Moderators to abstain from processing tasks only if and only if the task conflicted with their assigned values. Using GPT-4o, we conducted 30 iterations per (task, value) combination, resulting in 360 total annotations. In each iteration, a Moderator decides whether to accept or reject the given task.
@@ Appendix A Deliberation Ideals
Table 2. Translating ideals of deliberative democracy into instructions for LLMs. Starting from the taxonomy in Bächtiger et al. (bachtiger_deliberative_2018, ), two authors engaged in an iterative process where we first screened ideals fore relevance to AI agents and then translated ideals into LLM instructions.
@@ Appendix E Multilevel Logistic Regressions of Efficacy Studies
Table 3. Mixed effect logistic results from efficacy studies. Participants chose between Plurals or non-Plurals output. The outcome variable is choosing Plurals. The fixed effect intercept is the baseline odds of choosing Plurals output, accounting for random effects. Models 1-3 have a random intercept for participants. Model 4 collapses across studies, with a random intercept for participants crossed with a random intercept for studies.
@@
Constant 15.631 3.932 2.812 5.714
t = 5.559∗∗∗ t = 2.466∗∗ t = 2.518∗∗ t = 4.017∗∗∗
Random Intercept Variance (Person) 2.501 5.178 2.503 3.595
Random Intercept Variance (Study) 0.329
@@
Log Likelihood --93.969 --139.743 --109.423 --347.191
Akaike Inf. Crit. 191.937 283.486 222.846 700.383
Bayesian Inf. Crit. 199.345 290.894 229.443 714.436
@@ H.1. Plurals Code
52 task=critique_prompt,
@@
58 task=critique_prompt,
@@
64 task=critique_prompt,
@@
69 task=REVISE_PROMPT,

### v5

@@ Abstract.
Recent debates raised concerns that language models may favor certain viewpoints. But what if the solution is not to aim for a “view from nowhere” but rather to leverage different viewpoints? We introduce Plurals, a system and Python library for pluralistic AI deliberation. Plurals consists of Agents (LLMs, optionally with personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is a generator of simulated social ensembles. Plurals integrates with government datasets to create nationally representative personas, includes deliberation templates inspired by deliberative democracy, and allows users to customize both information-sharing structures and deliberation behavior within Structures. Six case studies demonstrate fidelity to theoretical constructs and efficacy. Three randomized experiments show simulated focus groups produced output resonant with an online sample of the relevant audiences (chosen over zero-shot generation in 75% of trials). Plurals is both a paradigm and a concrete system for pluralistic AI.
@@ 1. Introduction
As an alternative to “bias-free” models, we introduce a new pluralistic AI system (sorensen_roadmap_2024, ), Plurals, that can accomplish these tasks. It is a public-facing Python library (Figure 1 for system overview, Figure 2 for code snippets, see here [^1] https://github.com/josh-ashkinaze/plurals for library). Plurals consists of Agents (optionally integrated with government datasets for nationally representative personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is an end-to-end generator of customizable “simulated social ensembles”. We incorporate interaction templates inspired by democratic deliberation theory and integration with government datasets for nationally representative personas. For example, to create an Agent representing a male California resident, our system samples a statistically representative citizen from American National Election Studies, and then uses the citizen’s demographics and political stances as an LLM prompt. We draw on deliberative democracy theory, which emphasizes dialogue between different views (bachtiger_deliberative_2018, ; marti_pluralism_2017, ), as a blueprint. Our work builds on research in deliberation (bachtiger_deliberative_2018, ; morrell_listening_2018, ; brown_deliberation_2018, ; smith_mini-publics_2018, ; fraser_rethinking_1990, ; habermas_structural_1991, ; marti_pluralism_2017, ), pluralistic sociotechnical systems (argyle_out_2023, ; lee_webuildai_2019, ; gordon_jury_2022, ; zhang_policykit_2020, ), and multi-agent AI alignment approaches (vahidov_pluralistic_2004, ; irving_ai_2018, ; talebirad_multi-agent_2023, ; kim_can_2024, ; chan_chateval_2023, ). To our knowledge, Plurals is the first general-purpose toolkit for pluralistic, multi-agent interactions modeled after democratic deliberation.
@@ 1.1. Brief System Overview
Plurals allows users to create simulated social ensembles with Agents, Structures, and Moderators: Agents complete tasks within Structures, which define how information is shared between Agents. Moderators can summarize multi-agent communication. Each abstraction is highly customizable. Agents can use various LLMs and have system instructions set manually, through persona generation, or via American National Election Studies (ANES) integration. Structures vary in information-sharing, complexity, and randomness. For example, users can define custom networks of Agents in a few lines of code (Figure 2). The behavior of Agents within Structures (how they should combine information from other agents) can be tuned via combination instructions. Our package comes pre-populated with templates for personas, combination instructions, and moderators—drawing on democratic deliberation theory and prior work.
@@ 2.3. Multi-Agent Systems for AI Alignment
Like our system, several multi-agent systems are explicitly designed with the goal of alignment (irving_ai_2018, ; pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ). Broadly, these systems typically center interactions between agents or agent roles. For example, several projects have explored the role of AI alignment through debate( (irving_ai_2018, ; khan_debating_2024, )). Khan et al. (khan_debating_2024, ) propose a system where a weaker model picks an answer after seeing stronger models debate, and this process of debate leads to more truthful answers for the weaker model. Other multi-agent systems center agent roles (pang_self-alignment_2024, ; mangal_coalitions_2024, ; hu_agentscomerge_2024, ; zhu_role-based_2008, ; vahidov_pluralistic_2004, )—the idea being that agents playing distinct parts can aid human decision-makers (vahidov_pluralistic_2004, ).
To this body of research, we offer several contributions. More theoretically, our abstractions are specifically grounded in the theory and practice of deliberation. More practically, because our system has support for Agents, Structures, and Moderators, it effectively enables users to customize both information-sharing (as in AI debate literature) and Agent roles (as in the AI role literature). We extend the debate paradigm by allowing for arbitrary information structures. A back-and-forth debate is of course just one of many possible informational structures. Our system contributes to the role-based literature by integrating with ANES, enabling users to quickly draw up nationally representative roles. We also design around deliberation—the space in between roles and information-sharing. For example, users can ablate the role of an Agent (i.e.: their system instructions) and the combination instructions of an Agent. Finally, Plurals is a fully functioning Python package and not a one-off study. Hence, Plurals can operate as shared infrastructure. It makes multi-agent systems faster to set up and more accessible for researchers.
@@ 3.2. Modularity
Apart from the practical utility, this separation between Agents and Structures aligns with well-established social science frameworks. This conceptualization is most explicitly articulated in Structuration Theory by Anthony Giddens (giddens_constitution_1986, ), which explores the interplay between “agents” and larger “structures” agents exist in. Giddens aimed to transcend theories of behavior that centered exclusively on either individuals or societal structures. Similar distinctions appear across disciplines: individuals and environments in development psychology (radke-yarrow_individual_1991, ), person and situation in social psychology (furr_persons_2021, ), individual and field in sociology (grenfell_field_2012, ), and agent and environment in artificial intelligence (russell_artificial_1995, ). By using Agents and Structures as core abstractions [^2] We chose the terms “Agents” and “Structures” based on AI terminology and definitional precision. We adopt conventional AI terminology, where agents refer to autonomous entities capable of perceiving and acting within an environment (russell_artificial_1995, ). We opted for “structures” instead of “environment” to more accurately reflect our system. Structures specifically define information-sharing patterns and interaction protocols between agents, describing a more bounded space than “environment.” Unlike “environment,” which often implies indeterminacy or randomness, “structure” connotes intentional arrangement. The Cambridge Dictionary defines structure as “the way in which the parts of a system or object are arranged or organized, or a system arranged in this way” (dictionary_structure_2024, ). This definition aligns with a user-tuned abstraction., we create a modularity that resonates with different disciplines.
@@ 4.1.1. Component Description
(b) In a moderated ensemble, nationally representative Agents brainstorm ways to encourage recycling. Then a moderator with a persona inspired by divergent creativity literature (ashkinaze_how_2024, ) summarizes responses with custom combination instructions.
@@ 4.1.2. Implementation
Specific personas can be inputted by the user (e.g: “A graphic designer”) or drawn from American National Election Studies (ANES) [^4] Specifically, we are using the ANES pilot dataset from February 2024., as in Argyle et al. (argyle_out_2023, ). When using ANES, our system finds a real individual satisfying some criteria and then creates a persona based on the totality of this individual’s attributes. Sampling is always probability-weighted, so the probability of a citizen being simulated matches their national sample probability weight. Because ANES is nationally representative, the marginal distribution of Plurals-generated personas matches that of the general population. Code snippet 2(d) (top panel), shows initializing Agents based on specific criteria (e.g: California resident below the age of 40) using the query_str method, which searches ANES through a Pandas string [^5] For accessibility we have a helper function which prints a human-readable mapping of ANES variables.. For convenience, we also support an ideology method (ideology=’liberal’) and initializing randomly selected ANES citizens (persona=’random’, 2(a)). The latter can be used to quickly draw up nationally representative “citizen assemblies” (2(b)).
ANES is just one possible generator of data-driven personas, and in future iterations, we aim to provide additional persona-generation methods. We chose ANES as our initial dataset for the following reasons. First, it has been used in prior work—most notably, Argyle et al. (argyle_out_2023, ). Second, ANES has data on political ideologies, supporting the core motivation of this system—testing whether LLM outputs can be improved through pluralism. Third, ANES is updated more frequently than other nationally representative datasets like the U.S. census.
@@ Results
In the zero-shot condition, we set the system instructions of GPT-4o to “You are an expert copywriter for an ad agency” and the user prompt was “Come up with a specific product for a solar panel company that would resonate with conservatives. Be very specific. Answer in 50 words only.” In the Plurals condition, the Moderator had the same system instructions. However, that Moderator oversaw an ensemble of 10 simulated ANES conservatives (initialized using our ideology persona method and anes persona template) who were asked what features they personally would want in a solar panel company. The Moderator then came up with a 50-word solar panel idea after exposure to these simulated discussions. For 15 trials, we generated a solar panel company idea with zero-shot and Plurals.
@@
5.5. Efficacy: Liberals preferred homeless shelter ideas from a simulated focus group of liberals over zero-shot.
@@ 5.6. Moderation: Using Plurals for LLM Guardrails
We began with Abercrombie et al.’s (abercrombie_collaborative_2024, ) typology of AI, algorithmic, and automation harms. We selected two specific harms—environmental and physical harms. For each harm, we crafted three user prompts that would trigger concerns in one category but not the other (Appendix I), testing the Moderator’s ability to discriminate between tasks based on their specific value sets. We initialized Moderators with specific value sets using a CoT system prompt that incorporated Abercrombie et al.’s language around typology definitions (Appendix I), instructing Moderators to abstain from processing tasks if and only if the task conflicted with their assigned values. Using GPT-4o, we conducted 30 iterations per (task, value) combination, resulting in 360 total annotations. In each iteration, a Moderator decides whether to accept or reject the given task.
@@ Appendix A Deliberation Ideals
Table 2. Translating ideals of deliberative democracy into instructions for LLMs. Starting from the taxonomy in Bächtiger et al. (bachtiger_deliberative_2018, ), two authors engaged in an iterative process where we first screened ideals for relevance to AI agents and then translated ideals into LLM instructions.
@@ Appendix E Multilevel Logistic Regressions of Efficacy Studies
Table 3. Mixed effect logistic results from efficacy studies. Participants chose between Plurals or non-Plurals output. The outcome variable is choosing Plurals. Models 1-4 have a random intercept for participants. Model 4 collapses across studies. The fixed effect intercept represents the odds (exponentiated logit coefficient) of choosing our system for a typical participant.
@@
Constant 15.631 3.932 2.812 5.855
t = 5.559∗∗∗ t = 2.466∗∗ t = 2.518∗∗ t = 5.734∗∗∗
Random Intercept Variance (Person) 2.501 5.178 2.503 4.043
@@
Log Likelihood --93.969 --139.743 --109.423 --347.845
Akaike Inf. Crit. 191.937 283.486 222.846 699.690
Bayesian Inf. Crit. 199.345 290.894 229.443 709.059
@@ H.1. Plurals Code
52 task=feedback_prompt,
@@
58 task=feedback_prompt,
@@
64 task=feedback_prompt,
@@
69 task=feedback_prompt,